# Breaking Barriers: Do Reinforcement Post Training Gains Transfer To Unseen Domains?

### v2

@@ Abstract
Reinforcement post training (RPT) has recently shown promise in improving the reasoning abilities of large language models (LLMs). However, it remains unclear how well these improvements generalize to new domains, as prior work evaluates RPT models on data from the same domains used for fine-tuning. To understand the generalizability of RPT, we conduct two studies. (1) Observational: We compare a wide range of open-weight RPT models against their corresponding base models across multiple domains, including both seen and unseen domains in their fine-tuning data. (2) Interventional: we fine-tune LLMs with RPT on single domains and evaluate their performance across multiple domains. Both studies converge on the same conclusion that, although RPT brings substantial gains on tasks similar to the fine-tuning data, the gains generalize inconsistently and can vanish on domains with different reasoning patterns.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable performance across a wide range of tasks, from structured reasoning domains such as math [27, 13, 36, 70] and code [26, 11, 89, 7, 61, 43, 4], to general reasoning domains including legal [21], finance [86], and medical [30, 20]. Recent advances in reinforcement post training (RPT) [58] have produced dramatic improvements, rivaling top human performers in programming competitions and mathematics contests [39, 40, 57, 64, 87, 83, 14, 75, 66, 22, 37, 18, 74, 19, 51, 6]. This raises an important question: does RPT provide generalizable improvements, as broadly as those achieved through pretraining?
@@
Existing evaluation frameworks and RPT setups provide limited evidence to answer this question. To address it systematically, we design a two-stage investigation pipeline (Figure 1(a)).
@@
First, prior work evaluates RPT models within their fine-tuning domains [40, 39]. To overcome this limitation, we conduct an observational study in which we evaluate 14 recent open-weight RPT models with publicly disclosed fine-tuning data alongside their corresponding base models across a wide range of domains, including legal, financial, and medical benchmarks, spanning their seen and unseen domains. This study is designed to provide an initial view into the generalizability of RPT.
@@
Additionally, we notice that these RPT models, as a representative selection of existing open-weight models such as DeepSeek R1 [16] and RLVR [63], are fine-tuned on mixed domain data. The presence of such confounding factors makes it difficult to isolate and interpret the generalizability of RPT at a finer granularity. To strengthen our findings, we conduct an interventional study in which we fine-tune LLMs via RPT on math, coding, and general reasoning data and evaluate their performance on both in-domain and out-of-domain tasks. We illustrate our methodology in detail in Section 3.
@@
As we summarize in Figure 1(b), our findings show that gains from RPT on domains involving structured reasoning patterns (e.g., math, code) generalize well within and across structured domains, but fails to generalize to unstructured domains. In contrast, gains from RPT on unstructured domains (e.g., legal, finance) do not generalize well within unstructured domains, but show transferability to structured domains. We analyze these results comprehensively in Section 4.
@@
Our findings suggest that RPT models are most effective when the target task shares reasoning patterns with the RPT data. Consequently, while RPT remains a powerful method for improving LLMs’ performance, its benefits are largely limited to the domains represented in the fine-tuning data and do not generalize to a wide range of new, unseen domains.
@@ 2 Background
In this section, we introduce the motivation behind our study. We begin by demonstrating the strong performance of LLMs fine-tuned via RPT across various reasoning tasks, particularly in mathematics and coding. We then discuss the limitations of existing work in understanding the mechanisms and boundaries of RPT. We close by introducing the data-domain taxonomy grounds our investigations.
@@
RPT models demonstrate promising performance across a wide range of tasks. RPT models have achieved remarkable improvements on complex reasoning benchmarks. For example, Gemini 2.5 Pro [18] achieves over 90% accuracy on AIME 2024 [27], a math competition benchmark. Grok 3 Beta [74] solves roughly 80% of the tasks in LiveCodeBench v5 [26], a coding benchmark, while Claude 3.7 Sonnet [6] reaches around 85% accuracy on GPQA Diamond [54], a benchmark for graduate-level scientific reasoning.
@@
Out-of-domain generalizability of RPT models remains understudied. Despite impressive results, recent work has examined the limitations of RPT models and the opaque nature of their underlying reasoning capabilities [67, 84, 42, 81]. In particular, there is growing interest in the role of RPT data, especially the extent to which RPT algorithms rely on large, diverse corpora to achieve generalization [87].
@@
While RPT models benefit from training on diverse, mixed-domain data, this diversity makes it difficult to directly assess their generalizability to unseen domains. As a result, prior work evaluates RPT models on tasks within the same domains as their training data [67, 84, 42, 81]. However, many reasoning tasks remain underrepresented or entirely absent from existing training corpora. Exploring the generalizability of LLMs trained with RPT to such tasks is therefore essential for identifying the boundaries of their applicability in real-world, complex scenarios.
@@
Understanding RPT generalizability requires a systematic view of reasoning domains. Following the task suites defined in prior work [63], we focus on three major domains of interest: code, math, and general reasoning (Figure 1(b)). These domains are chosen to capture a broad spectrum of reasoning challenges commonly seen in language model evaluations. The general reasoning domain can be further divided into application-specific subdomains such as legal, finance, medical, etc. In this paper, we not only examine performance across the high-level domains, but also evaluate how reasoning patterns and generalization behaviors vary across subdomains.
@@
Among the three, we consider math and code domain tasks to follow structured reasoning patterns, where solutions follow deterministic logical steps and require precise syntax and formal semantics [63]. In contrast, tasks within the general reasoning domain require more flexible and context-sensitive reasoning, referred to here as unstructured reasoning. We define unstructured reasoning as problem-solving processes that do not adhere to a fixed sequence of logical operations and often lack a well-defined intermediate representation or symbolic grounding. Such tasks typically demand broader world knowledge, interpretive judgment, and the ability to handle ambiguity or incomplete information. For instance, legal and financial question answering may involve interpreting lengthy documents, extracting relevant information from loosely connected statements, or evaluating conflicting evidence.
@@ 3 Study Design
To address these research questions, we design a two-stage pipeline. First, we perform an observational study by evaluating 14 RPT models, each compared against its corresponding base model across a diverse set of benchmarks, spanning their seen and unseen domains. Because existing RPT models are typically trained with different configurations (e.g., different RL algorithms and hyperparameters) on multi-domain data, it is challenging to isolate the effect of RPT itself from the advantages brought by specific configuration or data.
To mitigate confounding factors, we further conduct an interventional study, where we fine-tunedd three RPT models from the same base model with the same configuration, each on a disjoint single-domain dataset. We then evaluate these trained models using the same benchmarks as in the observational analysis. In the rest of this section, we describe the evaluation settings and experimental setup for both studies.
@@
  • Math: For easier questions, we use GSM8K [13] and MATH-500 [36], while for more challenging problems, we select AIME 2024 [27] and AMC 2023 [70].
  • Code: We use easy coding problems, including MBPP [7] and HumanEval [11], and relatively challenging problems, including BigCodeBench [89], LiveCodeBench [26], USACO [61], and Codeforces [43]. To test programming language generalization, we also include the multi-language benchmark Polyglot [4].
  • General reasoning: We use high-quality benchmarks that are not mathematics nor programming problems for general reasoning, including PubMedQA [30] and MedQA [20] for medical reasoning, TabFact [12] for fact verification, LegalBench [21] for legal reasoning, and FinBench [86] for financial problem solving.
Evaluation Configurations. For all benchmarks, we use a consistent sets of generation hyperparameters across all models. The maximum response length is set to $\min\{16192,\,C\}$, where $C$ denotes the model’s context window. For each model, we run the small benchmarks (i.e., AMC 2023 and AIME 2024) 16 times, while executing all other benchmarks once.
For prompting, we apply each model’s default chat template and system prompt. For model pairs whose base models are pretrained-only (i.e., not instruction-tuned), we evaluate under two prompting strategies: (1) the default prompt, and (2) the official CoT+few-shot template used for evaluating Qwen2.5-Math [80], using $\min\{4,\,N\}$ shots, where $N$ is the maximum number of few-shot exemplars that can fit within the model’s context window. For each model (the reasoning model and its corresponding base model), we report the higher score across the two prompting strategies, ensuring that both models are evaluated under their most favorable prompting conditions. We evaluate each model on each benchmark on a single 24 GB RTX A5000.
Metrics. To assess whether RPT improves the accuracy performance within or across domains, we report the aggregated accuracy improvement $\Delta_{i,j}^{(\mathcal{D})}$ of an RPT model $i$ over its base model $j$ for a given domain $\mathcal{D}$:
@@
In addition, to ensure statistical significance in our findings, we applied the Cochran–Mantel–Haenszel (CMH) test [1], a statistical test for analyzing stratified categorical data. We treat each benchmark as an independent stratum—that is, a random sample of distinct downstream tasks. Given an RPT model $i$, a base model $j$, and a domain $\mathcal{D}$ of benchmarks, we calculate the common odds ratio estimate ($\theta_{i,j,\mathcal{D}}$) that estimates the correlation between the RPT process and the accuracy improvement on $\mathcal{D}$:
@@
An odds ratio greater than 1 indicates improvement due to RPT; a value less than 1 indicates a decrease in accuracy due to RPT. We evaluate the statistical significance under the null hypothesis $H_{0}:\theta_{i,j}^{(\mathcal{D})}=1$ against the alternative hypothesis $H_{1}:\theta_{i,j}^{(\mathcal{D})}\neq 1$, using the standard CHM test statistics,
@@
which follows a chi-squared distribution asymptotically with 1 degree of freedom.
@@
(1) DeepScaleR-1.5B-Preview [40] DeepSeek-R1-Distill-Qwen-1.5B [16] Math
(2) DeepCoder-1.5B-Preview [39] DeepSeek-R1-Distill-Qwen-1.5B [16] Code
(3) Skywork-o1-Open-Llama-3.1-8B [48] Llama-3.1-8B-Instruct [44] Code, Math
(4) Eurus-2-7B-PRIME [14, 83] Qwen2.5-Math-7B [80] Code, Math
(5) Absolute_Zero_Reasoner-Coder-3b [87] Qwen2.5-Coder-3B [25, 79] Code
(6) Absolute_Zero_Reasoner-Coder-7b [87] Qwen2.5-Coder-7B [25, 79] Code
(7) ZR1-1.5B [90] DeepSeek-R1-Distill-Qwen-1.5B [16] Code, Math
(8) Llama-3.1-Nemotron-Nano-8B-v1 [9] Llama-3.1-8B-Instruct [44] Instruction Following
(9) Thespis-Llama-3.1-8B [38] Meta-Llama-3.1-8B-Instruct-abliterated [46] Chat
(10) STILL-3-1.5B-preview [68, 28, 45] DeepSeek-R1-Distill-Qwen-1.5B [16] Math
(11) Arcee-Maestro-7B-Preview [3] DeepSeek-R1-Distill-Qwen-7B [16] Code, Math
(12) Fino1-8B [53] Llama-3.1-8B-Instruct [44] Finance
(13) OREAL-7B [41] OREAL-7B-SFT [41] Math
(14) Open-RS3 [15] DeepSeek-R1-Distill-Qwen-1.5B [16] Math
@@
  • Stage 2. We use o4-mini [50] to prefilter models potentially trained with RPT, based on their model card descriptions. This automatic filtering is followed by manual verification, resulting in 31 models that we confirm to be RPT models.
@@
Additionally, we include 2 recently released models, Absolute_Zero_Reasoner-Coder-3B and Absolute_Zero_Reasoner-Coder-7B [87], both fine-tunedd with limited RPT data and released on May 6th, 2025. These models demonstrate strong performance on math and code reasoning tasks and serve as representative cases for examining RPT generalizability.
We finalize our selection of 14 RPT models, with the details, including base models and RPT domains, presented in Table 1. For each RPT model and its corresponding base model, we compare performance across 16 benchmarks.
Interventional Analysis. To isolate the effect of RPT from other training configurations, including datasets, algorithms, and hyperparameters, we trained three RPT models based on DeepSeek-R1-Distill-Qwen-1.5B [16] on three disjoint datasets—math, code, and general reasoning—respectively. We curated these datasets based on existing datasets that leads to performant RPT models. Specifically,
  • Math: we uniformly sampled 40,000 problems from a combination of the math split of Eurus-2-RL [14], which originates from the NuminaMath-CoT dataset [33].
  • Code: we uniformly sampled 40,000 deduplicated problems from a combination of KodCode [78], DeepCoder-Preview [39], Apps [23], TACO [34], and the code split of Eurus-2-RL [14].
  • General Reasoning: we selected 40,000 high-quality, non-math, and non-code data from the multi-subject RLVR dataset [64]. To achieve that, we applied o3-mini [50] to exclude math-related, code-related, or fact-recall questions.
To ensure fair evaluation, we clean this data to ensure that it does not overlap with our evaluation set.
We applied consistent settings for all three RPT training processes. In terms of the RL algorithm, we applied Group Relative Policy Optimization (GRPO) with the same setting as DeepCoder [39]. In terms of hyperparameters, we trained each of the dataset for one epoch with a batch size of 64 and a context length of 8,192. To stabilize the training process, we used a learning rate of $10^{-6}$ and an entropy coefficient of 0. We fine-tuned the models on 8 80GB H100 GPUs.

### v3

# arXiv:2506.19733v3[cs.CL] 02 Mar 2026

@@ Abstract
Reinforcement post training (RPT) has recently shown promise in improving the reasoning abilities of large language models (LLMs). However, it remains unclear how well these improvements generalize to new domains, as prior work evaluates RPT models on data from the same domains used for post-training. To understand the generalizability of RPT, we conduct two studies with specific focus on Reinforcement Learning with Verifiable Rewards (RLVR). (1) Observational: we compare a wide range of open-weight RPT models against their corresponding base models across multiple domains, including both seen and unseen domains in their fine-tuning data. (2) Interventional: we fine-tune LLMs with RPT on single domains and evaluate their performance across multiple domains. Both studies converge on the same conclusion that, although RPT brings substantial gains on tasks similar to the fine-tuning data, the gains generalize inconsistently and can vanish on domains with different reasoning patterns.
@@ 1 Introduction
Large language models (LLMs) demonstrate strong capabilities across diverse reasoning settings. In mathematics and quantitative reasoning, recent models achieve near-expert performance on benchmarks including GSM8K, MATH, and AIME (Cobbe et al., 2021; Lightman et al., 2023; Jia, 2024; Tunstall and Jia, 2024). In code generation and program synthesis, LLMs similarly show substantial progress across challenging evaluations (Austin et al., 2021; Chen et al., 2021; Jain et al., 2024; Shi et al., 2024; Zhuo et al., 2025; MatrixStudio, 2024; Aider-AI, 2025). Beyond these structured domains, LLMs also perform well on knowledge-intensive reasoning spanning legal, financial, and biomedical tasks (Guha et al., 2023; Zhang et al., ; Jin et al., 2019; Griot et al., 2025). A growing body of work shows that reinforcement post-training (RPT) can yield large performance gains on these tasks. Recent RPT-enhanced models achieve dramatic improvements in competition-level mathematics and coding benchmarks, in some cases matching or exceeding top human competitors (Shao et al., 2024; Luo et al., 2025a; b; Su et al., 2025b; Zhao et al., 2025a; Yuan et al., 2024; Cui et al., 2025; Xie et al., 2025; Team, 2025a; He et al., 2025; Liu et al., 2025; Google DeepMind, 2025a; xAI, 2025; Granite Embedding Team, 2024; OpenAI et al., 2024; Anthropic, 2025a). This raises an important question: does RPT provide generalizable improvements, as broadly as those achieved through pretraining?
@@
Existing evaluation frameworks and RPT setups provide limited evidence to answer this question. To address it systematically, we design a two-stage investigation pipeline (Figure 1(a)), with a specific focus on Reinforcement Learning with Verifiable Rewards (RLVR), the dominant instantiation of RPT for verifiable reasoning tasks.
@@
First, prior work evaluates RPT models within their post-training domains (Luo et al., 2025b; a). To overcome this limitation, we conduct an observational study in which we evaluate 18 recent open-weight RPT models with publicly disclosed post-training data alongside their corresponding base models across a wide range of domains, including legal, financial, and medical benchmarks, spanning their seen and unseen domains. This study is designed to provide an initial view into the generalizability of RPT.
@@
Additionally, we notice that these RPT models, as a representative selection of existing open-weight models such as DeepSeek R1 (DeepSeek-AI et al., 2025), are post-trained on mixed domain data. The presence of such confounding factors makes it difficult to isolate and interpret the generalizability of RPT at a finer granularity. To strengthen our findings, we conduct an interventional study in which we post-train LLMs via RPT on math, coding, and knowledge-intensive reasoning data and evaluate their performance on both in-domain and out-of-domain tasks. We illustrate our methodology in detail in Section 3.
@@
As we summarize in Figure 1(b), our findings show that gains from RPT shows limited generalizabilty. Specifically, RPT gains on domains involving structured reasoning patterns (e.g., math, code) generalize well within and across structured domains, but fail to generalize to unstructured domains. In contrast, gains from RPT on unstructured domains (e.g., legal, finance) do not generalize well within unstructured domains, but show transferability to structured domains. We analyze these results comprehensively in Section 4.
@@
Our findings suggest that RPT models are most effective when the target task shares similar reasoning patterns with the RPT data. Consequently, while RPT remains a powerful method for improving LLMs’ performance, its benefits are largely limited to the domains represented in the post-training data and do not generalize to a wide range of new, unseen domains.
@@ 2 Background
In this section, we introduce the motivation of our study. We begin by demonstrating the strong performance of LLMs post-trained via RPT across various reasoning tasks, particularly in mathematics and coding. We then discuss the limitations of existing work in understanding the mechanisms and boundaries of RPT. We close by introducing the data-domain taxonomy that grounds our study.
@@
RPT models demonstrate promising performance across a wide range of tasks. RPT models have achieved remarkable improvements on complex reasoning benchmarks. For example, Gemini 3 Pro (Google DeepMind, 2025b) achieves 100% accuracy on AIME 2025 (OpenCompass Community, 2025), a math competition benchmark. Claude Opus 4.5 (Anthropic, 2025c) reaches 81% accuracy on SWE-bench Verified (Jimenez et al., 2024), a coding benchmark, while GPT-5.2-Pro (OpenAI, 2025a) reaches around 93.2% accuracy on GPQA Diamond (Rein et al., 2023), a benchmark for graduate-level scientific reasoning.
@@
Out-of-domain generalizability of RPT models remains understudied. Despite impressive results, recent work has examined the limitations of RPT models and the opaque nature of their underlying reasoning capabilities (Team, 2025b; Yue et al., 2025; Ma et al., 2025; Ye et al., 2025). In particular, there is growing interest in the role of RPT data, especially the extent to which RPT algorithms rely on large, diverse corpora to achieve generalization (Zhao et al., 2025a).
@@
While RPT models benefit from training on diverse, mixed-domain data, this diversity makes it difficult to directly assess their generalizability to unseen domains. As a result, prior work evaluates RPT models on tasks within the same domains as their training data (Team, 2025b; Yue et al., 2025; Ma et al., 2025; Ye et al., 2025). However, many reasoning tasks remain underrepresented or entirely absent from existing training corpora. Exploring the generalizability of LLMs trained with RPT to such tasks is therefore essential for identifying the boundaries of their applicability in real-world, complex scenarios.
@@
Understanding RPT generalizability requires a systematic view of reasoning domains. Following the task suites defined in prior work (Su et al., 2025a), we focus on three major domains of interest: code, math, and knowledge-intensive reasoning (Figure 1(b)). These domains are chosen to capture a broad spectrum of reasoning challenges commonly seen in language model evaluations. The knowledge-intensive reasoning domain can be further divided into application-specific subdomains such as legal, finance, medical, etc. We not only examine performance across the high-level domains, but also evaluate how reasoning patterns and generalization behaviors vary across subdomains.
@@
Among the three, we consider math and code domain tasks to follow structured reasoning patterns, where solutions follow deterministic logical steps and require precise syntax and formal semantics (Su et al., 2025a). In contrast, tasks within the knowledge-intensive reasoning domain require more flexible and context-sensitive reasoning, referred to here as unstructured reasoning. We define unstructured reasoning as problem-solving processes that do not adhere to a fixed sequence of logical operations and lack a well-defined intermediate representation or symbolic grounding. Such tasks demand broader world knowledge, interpretive judgment, and the ability to handle ambiguity or incomplete information. For instance, legal and financial question answering may involve interpreting lengthy documents, extracting relevant information from loosely connected statements, or evaluating conflicting evidence. We quantify the rationale of such divisions in Appendix B.
@@ 3 Study Design
  • (RQ4) Stability of generalization. Is the generalizability of RPT consistent across different hyperparameters, such as model size, training algorithm, and the number of RPT steps?
To address these research questions, we design a two-stage pipeline. First, we perform an observational study by evaluating 18 RPT models, each compared against its corresponding base model across a diverse set of benchmarks, spanning their seen and unseen domains. Because existing RPT models are typically trained with different configurations (e.g., different RL algorithms and hyperparameters) on multi-domain data, it is challenging to isolate the effect of RPT itself from the advantages brought by specific configuration or data.
To mitigate confounding factors, we further conduct an interventional study, where we post-trained three RPT models from the same base model with the same configuration, each on a disjoint single-domain dataset. We then evaluate these trained models using the same benchmarks as in the observational analysis. In the rest of this section, we describe our evaluation settings and experiment setups for both studies.
3.1 Evaluation Settings
@@
  • Math: For easier questions, we use GSM8K (Cobbe et al., 2021) and MATH-500 (Lightman et al., 2023), while for more challenging problems, we select AIME 2024 (Jia, 2024) and AMC 2023 (Tunstall and Jia, 2024).
  • Code: We use easy coding problems, including MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021), and relatively challenging problems, including BigCodeBench (Zhuo et al., 2025), LiveCodeBench (Jain et al., 2024), USACO (Shi et al., 2024), and Codeforces (MatrixStudio, 2024). To test programming language generalization, we also include the multi-language benchmark Polyglot (Aider-AI, 2025).
  • Knowledge-intensive reasoning: We use high-quality benchmarks that are not mathematics nor programming problems for knowledge-intensive reasoning, including PubMedQA (Jin et al., 2019) and MedQA (Griot et al., 2025) for medical reasoning, TabFact (Chen et al., 2019) for fact verification, LegalBench (Guha et al., 2023) for legal reasoning, and FinBench (Zhang et al., ) for financial problem solving.
Generation Configurations. For all benchmarks, we use a consistent set of generation hyperparameters across all models. The maximum response length is set to $\min\{16192,\,C\}$, where $C$ denotes the model’s context window. For each model, we run the small benchmarks (i.e., AMC 2023 and AIME 2024) 16 times, while executing all other benchmarks once. For prompting, we apply each model’s default chat template and system prompt.
Evaluation Metrics. To assess whether RPT improves the accuracy performance within or across domains, we report the aggregated accuracy improvement $\Delta_{i,j}^{(\mathcal{D})}$ of an RPT model $i$ over its base model $j$ for a given domain $\mathcal{D}$:
@@
In addition, to ensure statistical significance in our findings, we applied the Cochran–Mantel–Haenszel (CMH) test (Agresti, 2013), a statistical test for analyzing stratified categorical data. We treat each benchmark as an independent stratum—that is, a random sample of distinct downstream tasks. Given an RPT model $i$, a base model $j$, and a domain $\mathcal{D}$ of benchmarks, we calculate the common odds ratio estimate ($\theta_{i,j,\mathcal{D}}$) that estimates the correlation between the RPT process and the accuracy improvement on $\mathcal{D}$:
@@
An odds ratio greater than 1 indicates improvement from RPT; a value less than 1 indicates a decrease in accuracy from RPT. We evaluate the statistical significance under the null hypothesis $H_{0}:\theta_{i,j}^{(\mathcal{D})}=1$ against the alternative hypothesis $H_{1}:\theta_{i,j}^{(\mathcal{D})}\neq 1$, using the standard CHM test statistics,
@@
which follows a chi-squared distribution asymptotically with 1 degree of freedom. For all reported values, an asterisk (*) denotes statistical significance at $p<0.05$.
3.2 Experimental Settings
@@
(1) DeepScaleR-1.5B-Preview (Luo et al., 2025b) DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025) Math
(2) DeepCoder-1.5B-Preview (Luo et al., 2025a) DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025) Code
(3) Skywork-o1-Open-Llama-3.1-8B (Team, 2024) Llama-3.1-8B-Instruct (Meta AI, 2024b) Code, Math
(4) Eurus-2-7B-PRIME (Cui et al., 2025; Yuan et al., 2024) Eurus-2-7B-SFT (Cui et al., 2025; Yuan et al., 2024) Code, Math
(5) Absolute_Zero_Reasoner-Coder-3b (Zhao et al., 2025a) Qwen2.5-Coder-3B (Hui et al., 2024; Yang et al., 2024) Code
(6) Absolute_Zero_Reasoner-Coder-7b (Zhao et al., 2025a) Qwen2.5-Coder-7B (Hui et al., 2024; Yang et al., 2024) Code
(7) ZR1-1.5B (Zyphra, 2024) DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025) Code, Math
(8) Llama-3.1-Nemotron-Nano-8B-v1 (Bercovich et al., 2025) Llama-3.1-8B-Instruct (Meta AI, 2024b) Instruction Following
(9) Thespis-Llama-3.1-8B (Locutusque, 2024) Meta-Llama-3.1-8B-Instruct-abliterated (mlabonne, 2024) Chat
(10) STILL-3-1.5B-preview (Team, 2025d; Jiang et al., 2024; Min et al., 2024) DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025) Math
(11) Arcee-Maestro-7B-Preview (AI, 2024) DeepSeek-R1-Distill-Qwen-7B (DeepSeek-AI et al., 2025) Code, Math
(12) Fino1-8B (Qian et al., 2025) Llama-3.1-8B-Instruct (Meta AI, 2024b) Finance
(13) OREAL-7B (Lyu et al., 2025) OREAL-7B-SFT (Lyu et al., 2025) Math
(14) Open-RS3 (Dang and Ngo, 2025) DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025) Math
(15) DeepCoder-14B-Preview (Luo et al., 2025a) DeepSeek-R1-Distill-Qwen-14B (DeepSeek-AI et al., 2025) Code
(16) OREAL-32B (Lyu et al., 2025) OREAL-32B-SFT (Lyu et al., 2025) Math
(17) Fin-o1-14B (Qian et al., 2025) Qwen3-14B (Team, 2025c) Finance
(18)Absolute_Zero_Reasoner-Coder-14b (Zhao et al., 2025a) Qwen2.5-Coder-14B (Hui et al., 2024; Yang et al., 2024) Code
@@
  • Stage 2. We use o4-mini (OpenAI, 2025b) to prefilter models potentially trained with RPT, based on their model card descriptions. This automatic filtering is followed by manual verification, resulting in 31 models that we confirm to be RPT models.
@@
We evaluate Absolute_Zero_Reasoner-Coder-3B, post-trained with limited RPT data but demonstrating strong performance on math and code reasoning tasks. We view it as a representative case for examining RPT generalizability. We also evaluate the variants of all selected RPT models with more parameters, whenever such variants are available.
We finalize our selection of 18 RPT models, with the details, including base models and RPT domains, presented in Table 1. For each RPT model and its corresponding base model, we compare performance across 16 benchmarks.
Interventional Study. To isolate the effect of RPT from other training configurations, including datasets, algorithms, and hyperparameters, we trained three RPT models based on DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI et al., 2025) on three disjoint datasets, math, code, and knowledge-intensive reasoning, respectively. We curated training datasets based on existing datasets that led to performant RPT models and cleaned them to ensure that the training datasets do not overlap with our evaluation sets. We built the following datasets:
  • Math: we uniformly sampled 40,000 problems from a combination of the math split of Eurus-2-RL (Cui et al., 2025), which originates from the NuminaMath-CoT dataset (LI et al., 2024).
  • Code: we uniformly sampled 40,000 deduplicated problems from a combination of KodCode (Xu et al., 2025), DeepCoder-Preview (Luo et al., 2025a), Apps (Hendrycks et al., 2021), TACO (Li et al., 2023), and the code split of Eurus-2-RL (Cui et al., 2025).
  • Knowledge-intensive Reasoning: we selected 40,000 high-quality, non-math, and non-code data from the multi-subject RPT dataset (Su et al., 2025b). To achieve that, we applied o3-mini (OpenAI, 2025b) to exclude math-related, code-related, or fact-recall questions.
We applied consistent settings for all three RPT training processes. In terms of the RL algorithm, we applied Group Relative Policy Optimization (GRPO) with the same setting as DeepCoder (Luo et al., 2025a) , with the reward definitions and dynamics included in Appendix D. In terms of hyperparameters, we trained each dataset for one epoch with a batch size of 64 and a context length of 8,192. To stabilize the training process, we used a learning rate of $10^{-6}$ and an entropy coefficient of 0. We post-trained the models on 8 80GB H100 GPUs.
Furthermore, we conducted three additional experiments under alternative settings, to validate our conclusions about the generalizability of RPT,
  1. 1.
    We trained with DAPO (Yu et al., 2025), a state-of-the-art RL algorithm.
  2. 2.
    We extended the training process to 2 epochs and evaluated intermediate checkpoints.
  3. 3.
    We trained with Llama-3.2-3B-Instruct (Meta AI, 2024a) as the base model.